# KeyDiff: Key Similarity-Based KV Cache Eviction for Long-Context LLM Inference in Resource-Constrained Environments

### v3

@@ 1 Introduction
To enforce strict memory bounds throughout the prompt prefill and token generation inference phases, we adopt a block-wise inference strategy: the input prompt is divided into smaller blocks which are processed sequentially by the model, similar to [18, 1, 37]. After processing each block, we evict some of the cached KVs by according to an eviction policy that scores each KV, as illustrated in Figure 1. Unlike previous approaches that apply eviction after processing the entire prompt, this strategy satisfies memory constraints throughout the full inference process. However, we observe a degradation in accuracy when applying existing eviction methods in this setting (Table 1).
@@
  • Insight. We observe that lower pairwise cosine similarity among keys correlates with higher attention scores, suggesting its utility as a proxy for token importance. (Section 3.1)
  • Method. We introduce KeyDiff, an eviction strategy that selects keys based on their similarity to other cached entries without relying on attention scores or future tokens. (Section 3.2)
  • Theory. Through our analysis of key and query geometry we provide a theoretical understanding how/why KeyDiff works. We also show that KeyDiff solves an optimal subset selection problem that maximizes key diversity. (Section 3.3)
  • Performance. KeyDiff achieves $\leq 1.5\%$ and $\leq 0.04\%$ accuracy drop on LongBench with 6K and 8K cache budgets, respectively, outperforming state-of-the-art eviction methods across Llama and Qwen models (Section 4.2), and near non-evicting baseline performance for Deepseek-R1-Distill-Llama-8B on the Math-500 reasoning benchmark. (Section 4.3)
  • Efficiency. We observe up to $30\%$ end-to-end inference latency reduction using KeyDiff compared to existing KV cache eviction methods. (Section 4.5)
@@ 2.1 Transformers
where $O^{\operatorname{attn}}\in\mathbb{R}^{T\times d}$, and the causal attention mask $M$ is an upper triangular matrix with nonzero values of $\infty$- ∞.
@@ 2.2 KV Caching
When the $\operatorname{Attention}$ operator processes a new token $x_{T+1}$, it must also recompute the prior KV states for tokens $x_{0},\ldots,x_{T}$. This can be avoided by storing previously computed KVs in a KV cache $\mathcal{C}=(K,V)$ for later reuse and append the new KV corresponding $x_{T+1}$ to the cache. We can apply Equation 2 to an existing KV cache $\mathcal{C}$ as follows:
@@
where $k_{T+1}$, $q_{T+1}$, $v_{T+1}$ are the key, query, and value states of $x_{T+1}$, and $[X\|x_{T+1}]$ represents the concatenation of $x_{T+1}$ to an existing tensor $X$ along the time dimension, $M$ is the causal attention mask accounting for both the KV cache and $x_{T+1}$.
@@ 2.3 KV Cache Eviction Methods
Attention-based eviction methods prioritize KV pairs with higher attention scores to past tokens. This is problematic when applying block prompt processing: all input tokens are not simultaneously accessible within $\operatorname{Attention}$, only those in the current block and cache. This can result in an incorrect eviction decision. Additionally, attention-based eviction often requires explicitly materializing $A$, which can be resource intensive. We discuss the attention-based eviction policies further in Section 5.
@@ 2.4 KV Caching in Resource-Constrained Environments
As demonstrated in efficient LLM inference frameworks [1, 14, 18, 36], one solution is to apply $\pi_{N}$ more frequently by segmenting $X$ into non-overlapping blocks $X=[X_{0},X_{1},\ldots,X_{m-1}]$, where $X_{i}=[x_{Bi},\ldots,x_{B(i+1)-1}]$, $B$ is the block size, and $m=\lceil T/B\rceil$, and iteratively updating the cache by exploiting causality, applying Equation 4 in a block-wise fashion:
@@
where $k_{Bi:B(i+1)-1}$ and $v_{Bi:B(i+1)-1}$ are the keys and values selected from $X_{i}$ respectively, and $\mathcal{C}_{i}=(K_{i},V_{i})$ is the KV cache after processing the first $i$ prompt blocks. As in Equation 4, we concatenate the $B$ new KVs to the current cache, apply $\pi_{N}$ and update the cache in Equation 6.
@@ 3 Method
We demonstrate a negative correlation between attention scores and the cosine similarity among keys (Section 3.1) and leverage this observation to develop KeyDiff (Section 3.2), followed by a theoretical justification of KeyDiff (Section 3.3) and preliminary evidence of its efficacy (Section C.2).
@@ 3.1 Correlation of Attention Scores and Key Dissimilarity
To address the shortcomings of attention-based KV cache eviction in Section 2.4, we develop an alternative attention-free scoring metric that retains significant KVs across blocks while being resource efficient. We recall the “attention sink" phenomenon: LLMs often assign high attention weight to the first few tokens, regardless of the input [35, 29]; these highly weighted tokens are called sink tokens. However, the index of the sink tokens can vary across heads and layers and reside deeper in the sequence than the first few tokens. This observation motivates the following hypothesis: high attention scores can be determined by the intrinsic properties of the keys rather than by any particular combination of keys and queries.
@@
We evaluate our hypothesis by inspecting the cosine similarities between keys computed inside an attention block. We visualize the pairwise cosine similarities between keys along with the attention weights in two particular heads and layers in Figure 2. We observe that keys with lower cosine similarity with other keys exhibit higher relative attention scores regardless of the choice of query, such as the 4th and 15th keys in Figure 2(a), or the 1st key in Figure 2(b). Pairwise cosine similarity of keys is solely a function of the keys in the cache, which are independent of input queries; the surprising aspect of Figure 2 is the negative correlation with attention weights. These distinctive keys essentially recover the attention sink phenomenon [35].
@@ 3.2 KeyDiff
Based on the observation in Section 3.1, we propose KeyDiff, which evicts tokens from the KV cache based on key similarity. If the cache $\mathcal{C}$ has intermediate size $n$ and budget $N$ where $n>N$, $\pi^{\textsc{KeyDiff}{}}_{N}$ is defined as:
@@
Unlike attention-based eviction policies, KeyDiff does not require access to the attention weights $A$, facilitating optimized attention kernels that do not materialize $A$ such as FlashAttention [8]. However, computing the pairwise cosine similarities runs in $\mathcal{O}(n^{2})$ time. Fortunately, we can compute the score of each token in Equation 7 in $\mathcal{O}(n)$ as follows:
@@
where $\mu(\hat{K})=\frac{1}{n}\sum_{i=1}^{n}\hat{k}_{i}$ and $\hat{k}_{i}=\frac{k_{i}}{||k_{i}||}$. We refer to $\mu(\hat{K})$ as the anchor vector. We show this formulation retains the same KVs of Equation 7 under a mild condition. (see Section C.2). Our experimentation has shown that the anchor vector $\mu(\hat{K})$ can be replaced with $\mu(K)$ without losing accuracy (see Table 14). We evaluate the efficient KeyDiff described in Figure 3 using unnormalized keys $k$ in all subsequent sections. Figures 5, 8, 9 and 10 visualize the keys retained and evicted by sink attention [35], TOVA [24] and KeyDiff via PCA. KeyDiff retains more varied keys. A full complexity and FLOP analysis can be found in Appendices B and B.1
@@
In tasks such as reasoning and coding, where the most recent tokens are often important, we can augment KeyDiff and its efficient variant to use a percentage of the cache budget for a sliding window [6], which we call KeyDiff with sliding window. This extension introduces no complexity or memory overhead and we observe better results on certain tasks than vanilla KeyDiff (Tables 13 and E).
@@ 3.3 Why KeyDiff Works: A Theoretical Perspective
To solidify the theoretical foundation of KeyDiff and show that KeyDiff ultimately selects keys most aligned with queries, we prove the following two results. We first validate the relationship between cosine similarity and attention scores observed in Figure 2 by bounding the attention score of a new incoming key $k^{\star}$ in terms cosine similarity with a fixed query $q$:
@@
Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $||k_{i}||_{2}^{2}<M,\;\forall\;i$. Without loss of generality suppose $||q||=1$ and assume ${k^{*}}$ is a key not in $\{k_{i}\}_{i=1}^{n}$ with $||{k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$. Then, for $n\to\infty$,
@@
By combining Lemmas 3.1 and 3.2, we establish a relationship between the attention weight $w$ and the KeyDiff score $\mathrm{CosSim}(\bar{k},k^{*})$. As $\mathrm{CosSim}(\bar{k},q)$ decreases and $\mathrm{CosSim}(k^{*},q)$ increases (along with the attention weight $w$), then $\mathrm{CosSim}(\bar{k},k^{*})$ tends to $-1$: this means KeyDiff selects distinct keys most aligned with $q$. We visualize this in Figure 4 with a PCA embedding of keys and queries from a single head of Llama 3.2 3B, highlighting the relationship between top scoring keys via KeyDiff, the anchor vector and queries. Similar trends are found from the other layers and heads as shown in Figure 11. The proofs of Lemmas 3.1 and 3.2 are in Section C.3, along with empirical motivation for the chosen assumptions.
@@
Figure 5: (a, b, and c) PCA Visualizations in two dimensions of a key cache managed with Sink, TOVA, and KeyDiff. Retained tokens are blue, while evicted tokens are orange. Keys are taken from layer $5$ and head $3$ of Llama3.2-3B-Instruct, and generated using the NarrativeQA dataset. (d) PCA visualization of the retained keys for each KV cache eviction method.
@@ 4 Experiments
  • Needle-In-a-Haystack. KeyDiff outperforms competing eviction policies on the Needle-In-A-Haystack benchmark (Section 4.1).
  • LongBench. KeyDiff outperforms competing eviction policies with block size $B=128$ on LongBench, achieving an 1.5% accuracy drop with a 6K cache budget ($\sim$33compression rate) and $\leq$ .04% with a 8k cache budget ($\sim$23compression rate) with Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct (Section 4.2).
  • Reasoning. KeyDiff performs competitively on the Math-500 reasoning benchmark with other eviction methods using the DeepSeek-R1-Distill-Qwen-7B and Llama-8B, and shows near eviction-free baseline performance when augmented with a sliding window (Section 4.3) for DeepSeek-R1-Distill-Llama-8B.
  • Ablation Study. We perform an ablation study on the main parameters of KeyDiff and show that utilizing negative cosine similarity as the eviction criteria and the mean of cached keys as the anchor vector performs best. (Section 4.4).
  • Efficiency. We compare the end-to-end inference latency of KeyDiff, [20] and [24] and observe a 30% latency improvement with KeyDiff (Section 4.5).
@@
We apply several cache eviction methods to several decoder-only transformer-based language models, including Llama $3.1$-$8$B-Instruct [10], Llama $3.2$-$3$B-Instruct [10], and Qwen $2.5$-$3$B/7B-Instruct [38]. We evaluate these models using H2O [43], TOVA [24], SnapKV [20], and StreamingLLM [35], (or “sink attention") cache eviction policies, along with the eviction-free model as a baseline. We simulate a resource constrained environment by processing prompts and generating responses using Equation 6, with a block size of $B=128$ for prompt processing and $B=1$ for token generation using greedy decoding for all experiments. We denote the cache budgets of 2048, 4096, 6144 and 8192 as 2K, 4K, 6K and 8K, respectively.
@@
Figure 6: Accuracy across document length and needle depth for needle in a haystack test. Cache size is $6$K with $B=128$.
@@ 4.1 Needle In a Haystack
To compare the impact of various cache eviction policies on fact retrieval, we conduct the “Needle In a Haystack" test [21, 16]. This test embeds specific information (“needle") at different points within a body of unrelated text (“haystack"); finding and retaining the needle is challenging for eviction policies, which can’t know what information must be retained during block prompt processing. The results are shown in Figure 6 and Figure 19, where we show the recall accuracy of Llama3.23.23.23.2-$3$B-Instruct across different document lengths (x-axis) and needle depths (y-axis) with a cache size of $6$K. KeyDiff performs similarly to TOVA, SnapKV and sink attention for shorter documents and outperforms all three methods as the document length increases.
Table 1: Llama-3.1-8B/3.2-3B-Instruct LongBench results with $B=128$ (Higher is better). We highlight the best and second best methods within a given budget with bold and underline. We omit Chinese dataset results and other model results due to space limit. The full evaluation results are in Table 10. †: A subset of samples (183/200) were evaluated due to OOM errors.
@@ 4.2 LongBench
LongBench [4] is a bilingual, multi-task benchmark suite for LLMs, providing a comprehensive stress test for long prompt inputs. LongBench is useful for evaluating cache eviction methods in a resource constrained environments with a fixed memory budget: 51% of prompts are longer than the largest KV cache size of 8K. For cache budgets of 6k and 8k tokens, prompts in LongBench are compressed by 33% and 23% respectively on average, (see Section F.3 for more detail.)
Table 1 summarizes the evaluation results of Llama 3.1-8B-Instruct and Llama 3.2-3B-Instruct on the English subset of LongBench with 2K, 4K, 6K, and 8K cache budgets using various eviction policies with block prompt processing enabled with $B=128$. As shown in Table 1, KeyDiff outperforms other eviction strategies across most tasks, even demonstrating better performance with smaller cache budgets. KeyDiff shows significant a improvement on the PassageRetrieval-en (PR-en) dataset, which tests whether long-term dependencies within a long prompt can be correctly recognized [4], while achieving near full-context model performance even with the smallest budget. Adding a sliding window to KeyDiff improves coding task performance (Table 13). We observed similar trends in the full LongBench task suite as shown in Table 10 and in the additional results in Appendix F.
@@
We present the full evaluation results in Table 10 and more complete comparisons on LongBench in Appendix F, such as: standard prompt processing with a single large block (i.e. $B=\infty$) in Table 9; eviction method performance with Qwen 2.5-3B/7B-Instruct in Table 11; performance behavior with block sizes $B=[64,256]$ in Table 12; and performance on KeyDiff combined with a sliding window as described in Section 3.2. We also compare against the $L_{2}$-norm minimizing eviction method of [9] in Table 10.

### v4

@@ 1 Introduction
To enforce strict memory bounds throughout the prompt prefill and token generation inference phases, we adopt a block-wise inference strategy: the input prompt is divided into smaller blocks which are processed sequentially by the model, similar to [18, 1, 37]. After processing each block, we evict some of the cached KVs by according to an eviction policy that scores each KV, as illustrated in Figure˜1. Unlike previous approaches that apply eviction after processing the entire prompt, this strategy satisfies memory constraints throughout the full inference process. However, we observe a degradation in accuracy when applying existing eviction methods in this setting (Table˜1).
@@
  • Insight. We observe that lower pairwise cosine similarity among keys correlates with higher attention scores, suggesting its utility as a proxy for token importance. (Section˜3.1)
  • Method. We introduce KeyDiff, an eviction strategy that selects keys based on their similarity to other cached entries without relying on attention scores or future tokens. (Section˜3.2)
  • Theory. Through our analysis of key and query geometry we provide a theoretical understanding how/why KeyDiff works. We also show that KeyDiff solves an optimal subset selection problem that maximizes key diversity. (Section˜3.3)
  • Performance. KeyDiff achieves $\leq 1.5\%$ and $\leq 0.04\%$ accuracy drop on LongBench with 6K and 8K cache budgets, respectively, outperforming state-of-the-art eviction methods across Llama and Qwen models (Section˜4.2), and near non-evicting baseline performance for Deepseek-R1-Distill-Llama-8B on the Math-500 reasoning benchmark. (Section˜4.3)
  • Efficiency. We observe up to $30\%$ end-to-end inference latency reduction using KeyDiff compared to existing KV cache eviction methods. (Section˜4.5)
@@ 2.1 Transformers
where $O^{\operatorname{attn}}\in\mathbb{R}^{T\times d}$, and the causal attention mask $M$ is an upper triangular matrix with nonzero values of -$\infty$.
@@ 2.2 KV Caching
When the $\operatorname{Attention}$ operator processes a new token $x_{T+1}$, it must also recompute the prior KV states for tokens $x_{0},\ldots,x_{T}$. This can be avoided by storing previously computed KVs in a KV cache $\mathcal{C}=(K,V)$ for later reuse and append the new KV corresponding $x_{T+1}$ to the cache. We can apply Equation˜2 to an existing KV cache $\mathcal{C}$ as follows:
@@
where $k_{T+1}$, $q_{T+1}$, $v_{T+1}$ are the key, query, and value states of $x_{T+1}$, and $X\|x_{T+1}]$ represents the concatenation of $x_{T+1}$ to an existing tensor $X$ along the time dimension, $M$ is the causal attention mask accounting for both the KV cache and $x_{T+1}$.
@@ 2.3 KV Cache Eviction Methods
Attention-based eviction methods prioritize KV pairs with higher attention scores to past tokens. This is problematic when applying block prompt processing: all input tokens are not simultaneously accessible within $\operatorname{Attention}$, only those in the current block and cache. This can result in an incorrect eviction decision. Additionally, attention-based eviction often requires explicitly materializing $A$, which can be resource intensive. We discuss the attention-based eviction policies further in Section˜5.
@@ 2.4 KV Caching in Resource-Constrained Environments
As demonstrated in efficient LLM inference frameworks [1, 14, 18, 36], one solution is to apply $\pi_{N}$ more frequently by segmenting $X$ into non-overlapping blocks $X=[X_{0},X_{1},\ldots,X_{m-1}]$, where $X_{i}=[x_{Bi},\ldots,x_{B(i+1)-1}]$, $B$ is the block size, and $m=\lceil T/B\rceil$, and iteratively updating the cache by exploiting causality, applying Equation˜4 in a block-wise fashion:
@@
where $k_{Bi:B(i+1)-1}$ and $v_{Bi:B(i+1)-1}$ are the keys and values selected from $X_{i}$ respectively, and $\mathcal{C}_{i}=(K_{i},V_{i})$ is the KV cache after processing the first $i$ prompt blocks. As in Equation˜4, we concatenate the $B$ new KVs to the current cache, apply $\pi_{N}$ and update the cache in Equation˜6.
@@ 3 Method
We demonstrate a negative correlation between attention scores and the cosine similarity among keys (Section˜3.1) and leverage this observation to develop KeyDiff (Section˜3.2), followed by a theoretical justification of KeyDiff (Section˜3.3) and preliminary evidence of its efficacy (Section˜C.2).
@@ 3.1 Correlation of Attention Scores and Key Dissimilarity
To address the shortcomings of attention-based KV cache eviction in Section˜2.4, we develop an alternative attention-free scoring metric that retains significant KVs across blocks while being resource efficient. We recall the “attention sink" phenomenon: LLMs often assign high attention weight to the first few tokens, regardless of the input [35, 29]; these highly weighted tokens are called sink tokens. However, the index of the sink tokens can vary across heads and layers and reside deeper in the sequence than the first few tokens. This observation motivates the following hypothesis: high attention scores can be determined by the intrinsic properties of the keys rather than by any particular combination of keys and queries.
@@
We evaluate our hypothesis by inspecting the cosine similarities between keys computed inside an attention block. We visualize the pairwise cosine similarities between keys along with the attention weights in two particular heads and layers in Figure˜2. We observe that keys with lower cosine similarity with other keys exhibit higher relative attention scores regardless of the choice of query, such as the 4th and 15th keys in Figure˜2(a), or the 1st key in Figure˜2(b). Pairwise cosine similarity of keys is solely a function of the keys in the cache, which are independent of input queries; the surprising aspect of Figure˜2 is the negative correlation with attention weights. These distinctive keys essentially recover the attention sink phenomenon [35].
@@ 3.2 KeyDiff
Based on the observation in Section˜3.1, we propose KeyDiff, which evicts tokens from the KV cache based on key similarity. If the cache $\mathcal{C}$ has intermediate size $n$ and budget $N$ where $n>N$, $\pi^{\textsc{KeyDiff}{}}_{N}$ is defined as:
@@
Unlike attention-based eviction policies, KeyDiff does not require access to the attention weights $A$, facilitating optimized attention kernels that do not materialize $A$ such as FlashAttention [8]. However, computing the pairwise cosine similarities runs in $\mathcal{O}(n^{2})$ time. Fortunately, we can compute the score of each token in Equation˜7 in $\mathcal{O}(n)$ as follows:
@@
[…]
@@
In tasks such as reasoning and coding, where the most recent tokens are often important, we can augment KeyDiff and its efficient variant to use a percentage of the cache budget for a sliding window [6], which we call KeyDiff with sliding window. This extension introduces no complexity or memory overhead and we observe better results on certain tasks than vanilla KeyDiff (Tables˜14 and E).
@@ 3.3 Why KeyDiff Works: A Theoretical Perspective
To solidify the theoretical foundation of KeyDiff and show that KeyDiff ultimately selects keys most aligned with queries, we prove the following two results. We first validate the relationship between cosine similarity and attention scores observed in Figure˜2 by bounding the attention score of a new incoming key $k^{\star}$ in terms cosine similarity with a fixed query $q$:
@@
Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $k_{i}||_{2}^{2}<M,\;\forall\;i$. Without loss of generality suppose q=1||q||=1 and assume $k^{*}$} is a key not in $\{k_{i}\}_{i=1}^{n}$ with ${k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$. Then, for $n\to\infty$,
@@
By combining Lemmas˜3.1 and 3.2, we establish a relationship between the attention weight $w$ and the KeyDiff score $\mathrm{CosSim}(\bar{k},k^{*})$. As $\mathrm{CosSim}(\bar{k},q)$ decreases and $\mathrm{CosSim}(k^{*},q)$ increases (along with the attention weight $w$), then $\mathrm{CosSim}(\bar{k},k^{*})$ tends to 1-1: this means KeyDiff selects distinct keys most aligned with $q$. We visualize this in Figure˜4 with a PCA embedding of keys and queries from a single head of Llama 3.2 3B, highlighting the relationship between top scoring keys via KeyDiff, the anchor vector and queries. Similar trends are found from the other layers and heads as shown in Figure˜11. The proofs of Lemmas˜3.1 and 3.2 are in Section˜C.3, along with empirical motivation for the chosen assumptions.
@@
Figure 5: (a, b, and c) PCA Visualizations in two dimensions of a key cache managed with Sink, TOVA, and KeyDiff. Retained tokens are blue, while evicted tokens are orange. Keys are taken from layer 55 and head $3$ of Llama3.2-3B-Instruct, and generated using the NarrativeQA dataset. (d) PCA visualization of the retained keys for each KV cache eviction method.
@@ 4 Experiments
  • Needle-In-a-Haystack. KeyDiff outperforms competing eviction policies on the Needle-In-A-Haystack benchmark (Section˜4.1).
  • LongBench. KeyDiff outperforms competing eviction policies with block size $B=128$ on LongBench, achieving an 1.5% accuracy drop with a 6K cache budget ($\sim$33% compression rate) and $\leq$ .04% with a 8k cache budget ($\sim$23% compression rate) with Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct (Section˜4.2).
  • Reasoning. KeyDiff performs competitively on the Math-500 reasoning benchmark with other eviction methods using the DeepSeek-R1-Distill-Qwen-7B and Llama-8B, and shows near eviction-free baseline performance when augmented with a sliding window (Section˜4.3) for DeepSeek-R1-Distill-Llama-8B.
  • Ablation Study. We perform an ablation study on the main parameters of KeyDiff and show that utilizing negative cosine similarity as the eviction criteria and the mean of cached keys as the anchor vector performs best. (Section˜4.4).
  • Efficiency. We compare the end-to-end inference latency of KeyDiff, [20] and [24] and observe a 30% latency improvement with KeyDiff (Section˜4.5).
@@
We apply several cache eviction methods to several decoder-only transformer-based language models, including Llama 3.13.1-88B-Instruct [10], Llama 3.23.2-33B-Instruct [10], and Qwen 2.52.5-33B/7B-Instruct [38]. We evaluate these models using H2O [43], TOVA [24], SnapKV [20], and StreamingLLM [35], (or “sink attention") cache eviction policies, along with the eviction-free model as a baseline. We simulate a resource constrained environment by processing prompts and generating responses using Equation˜6, with a block size of $B=128$ for prompt processing and $B=1$ for token generation using greedy decoding for all experiments. We denote the cache budgets of 2048, 4096, 6144 and 8192 as 2K, 4K, 6K and 8K, respectively.
@@
Figure 6: Accuracy across document length and needle depth for needle in a haystack test. Cache size is 66K with $B=128$.
@@ 4.1 Needle In a Haystack
To compare the impact of various cache eviction policies on fact retrieval, we conduct the “Needle In a Haystack" test [21, 16]. This test embeds specific information (“needle") at different points within a body of unrelated text (“haystack"); finding and retaining the needle is challenging for eviction policies, which can’t know what information must be retained during block prompt processing. The results are shown in Figure˜6 and Figure˜19, where we show the recall accuracy of Llama3.23.2-33B-Instruct across different document lengths (x-axis) and needle depths (y-axis) with a cache size of 66K. KeyDiff performs similarly to TOVA, SnapKV and sink attention for shorter documents and outperforms all three methods as the document length increases.
Table 1: Llama-3.1-8B/3.2-3B-Instruct LongBench results with $B=128$ (Higher is better). We highlight the best and second best methods within a given budget with bold and underline. We omit Chinese dataset results and other model results due to space limit. The full evaluation results are in Table˜11. †: A subset of samples (183/200) were evaluated due to OOM errors.
@@ 4.2 LongBench
LongBench [4] is a bilingual, multi-task benchmark suite for LLMs, providing a comprehensive stress test for long prompt inputs. LongBench is useful for evaluating cache eviction methods in a resource constrained environments with a fixed memory budget: 51% of prompts are longer than the largest KV cache size of 8K. For cache budgets of 6k and 8k tokens, prompts in LongBench are compressed by 33% and 23% respectively on average, (see Section˜F.3 for more detail.)
Table˜1 summarizes the evaluation results of Llama 3.1-8B-Instruct and Llama 3.2-3B-Instruct on the English subset of LongBench with 2K, 4K, 6K, and 8K cache budgets using various eviction policies with block prompt processing enabled with $B=128$. As shown in Table˜1, KeyDiff outperforms other eviction strategies across most tasks, even demonstrating better performance with smaller cache budgets. KeyDiff shows significant a improvement on the PassageRetrieval-en (PR-en) dataset, which tests whether long-term dependencies within a long prompt can be correctly recognized [4], while achieving near full-context model performance even with the smallest budget. Adding a sliding window to KeyDiff improves coding task performance (Table˜14). We observed similar trends in the full LongBench task suite as shown in Table˜11 and in the additional results in Appendix˜F.
@@
We present the full evaluation results in Table˜11 and more complete comparisons on LongBench in Appendix˜F, such as: standard prompt processing with a single large block (i.e. $B=\infty$) in Table˜10; eviction method performance with Qwen 2.5-3B/7B-Instruct in Table˜12; performance behavior with block sizes $B=[64,256]$ in Table˜13; and performance on KeyDiff combined with a sliding window as described in Section˜3.2. We also compare against the $L_{2}$-norm minimizing eviction method of [9] in Table˜11.